# Conformalized Strategy-Proof Auctions

### v4

@@ 3 Proposed Method
Definition 1 (Statistically Strategy-Proof Auctions).
@@
In words, an auction outcome that satisfies Definition 1 is guaranteed to have low maximum regret with high probability, where the maximal regret value $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$ and the confidence level $1-\alpha$ are pre-defined by the auctioneer. Since we can control the maximal regret value with high probability, we refer to it as a statistically strategy-proof auction mechanism. Notably, this statistical notion of strategy-proofness does not imply zero regret at test time.
@@ 3.1 Regret Estimation
We propose two different approaches to design the regret estimation model ${r}_{{\hat{\theta}}_{r}}(b)$. In the shared backbone setup illustrated in Figure 3, the regret model is integrated into the existing allocation and payment backbone, providing a unified structure. The advantage of this approach is that it leverages the backbone features of the auction model to better predict the regret. However, this setup requires access to the auction model parameters and modifies the training procedure. In the second black-box approach, the auction model is treated as a black-box, avoiding the need to access the auction model’s internal parameters. This design choice is visualized in Figure 4, where the regret estimation model is a standalone neural network.
Figure 3: Shared backbone regret estimation.
Figure 4: Black-box auction model regret estimation.
@@ 3.2 Model Calibration
In essence, we show how to form a corrected threshold on the output ${r}^{\textup{max}}_{\hat{\theta}_{r}}(b_{\text{test}})$, ensuring the regret of the auction outcome obtained from the auction model does not exceed the requested maximal regret level $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$ with probability $1-\alpha$, resulting in a calibrated auction acceptance rule function. Given the auction model ${f}_{\hat{\theta}}$, a calibration set $\mathcal{D}_{\textup{cal}}$, a requested regret level $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$, and a control level $\alpha$, we define the auction acceptance rule function $\hat{\mathcal{T}}_{\alpha}(b_{\textup{test}},\operatorname{\ell_{\operatorname{\textbf{rgt}}}})$ for a test bid vector $b_{\text{test}}$ as follows:
$\hat{\mathcal{T}}_{\alpha}(b_{\text{test}},\operatorname{\ell_{\operatorname{\textbf{rgt}}}})=\begin{cases}{f}_{\hat{\theta}}(b_{\text{test}}),&{r}^{\textup{max}}_{\hat{\theta}_{r}}(b_{\text{test}})\leq\operatorname{\ell_{\operatorname{\textbf{rgt}}}}-\hat{q}_{\alpha},\\ \;\;\;\emptyset,&\text{otherwise}.\end{cases}$ (9)
The function $\hat{\mathcal{T}}_{\alpha}(b_{\text{test}},\operatorname{\ell_{\operatorname{\textbf{rgt}}}})$ from above is determined by the estimated maximum regret ${r}^{\textup{max}}_{\hat{\theta}_{r}}(b_{\text{test}})$ and the calibration threshold $\hat{q}_{\alpha}$, which is derived using the calibration set $\mathcal{D}_{\textup{cal}}$. The output of the function $\hat{\mathcal{T}}_{\alpha}(b_{\text{test}},\operatorname{\ell_{\operatorname{\textbf{rgt}}}})$ is the prediction of the auction model ${f}_{\hat{\theta}}({b_{\text{test}}})$ if ${r}^{\textup{max}}_{\hat{\theta}_{r}}(b_{\text{test}})$ is lower than the difference between the requested regret level $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$, and the correction factor $\hat{q}_{\alpha}$ derived from our calibration method. Otherwise, the function $\hat{\mathcal{T}}_{\alpha}$ yields an empty set $\emptyset$, indicating that the predicted auction does not meet the desired regret level. This acceptance rule ensures that the auction mechanism only operates within the predefined regret bounds.
@@
1:Input: ${f}_{\hat{\theta}}$ – pre-trained auction model; $r_{\hat{\theta}_{r}}$ – pre-trained regret model; $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$ – max regret level; $(1-\alpha)$ – requested confidence level; $\mathcal{D}_{\textup{cal}}$ – holdout calibration set.
@@
4:     Compute the estimated maximum regret ${r}_{\hat{\theta}_{r}}^{\textup{max}}(b_{k})$.
5:     Compute the actual maximum regret $\operatorname{\textbf{rgt}}_{\hat{\theta}}^{\textup{max}}(v_{k};b_{k})$.
6:     Calculate the score $e_{k}=\operatorname{\textbf{rgt}}_{\hat{\theta}}^{\textup{max}}(v_{k};b_{k})-{r}_{\hat{\theta}_{r}}^{\textup{max}}(b_{k})$.
@@
Table 1: Summary of the average revenue and empirical regret, with their standard deviations in parentheses, for both RegretNet and ConfRegretNet across $2\times 2$, $2\times 3$, and $3\times 5$ auction settings. The test and the calibration sets are of size 10,000 examples. The ‘Requested Max Regret Level‘ column indicates the requested regret used in ConfRegretNet to control the regret at $99\%$ confidence ($\alpha=0.01$). The ‘Accepted Auctions‘ column shows the percentage of accepted auctions by ConfRegretNet based on the regret threshold. the ‘Max Regret‘ column refers the max regret obtained at test time.
Auction Model Auctio Setup Bidders × Items Revenue Empirical Regret Requested Max Regret Max Regret Acceptance Auctions
RegretNet $2\times 2$ 0.93 (0.33) 0.007 (0.006) - 0.039 -
ConfRegretNet [ours] $2\times 2$ 0.92 (0.32) 0.004 (0.003) 0.025 0.020 89.23%
RegretNet $2\times 3$ 1.41 (0.38) 0.011 (0.014) - 0.164 -
ConfRegretNet [ours] $2\times 3$ 1.38 (0.39) 0.007 (0.005) 0.035 0.032 94.4%
RegretNet $3\times 5$ 2.93 (0.40) 0.023 (0.012) - 0.121 -
ConfRegretNet [ours] $3\times 5$ 2.67 (0.38) 0.009 (0.007) 0.055 0.055 93.4%
We now prove that any test auction accepted by the rule $\hat{\mathcal{T}}_{\alpha}$ satisfies Definition 1. Simply put, this result guarentees that by setting $\alpha=0.01$ and specifying a maximal regret level of $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}=0.025$, the actual regret of new auctions accepted by $\hat{\mathcal{T}}_{\alpha}$ will not exceed 0.025 with probability 99%.
@@
Suppose the calibration set $\mathcal{D}_{\textup{cal}}=\{(b_{k},v_{k})\}_{k=1}^{K}$ and a new auction $(b_{\textup{test}},v_{\textup{test}})$ are i.i.d samples from $P_{b,v}$. Let $\hat{\mathcal{T}}_{\alpha}$ be the acceptance rule of a given auction model ${f}_{\hat{\theta}}$ and regret estimator $r_{\hat{\theta}_{r}}$, obtained from Algorithm 1. Then, the auction ${f}_{{\hat{\theta}}}(b_{\textup{test}})$ accepted by the rule $\hat{\mathcal{T}}_{\alpha}$ in (9) satisfies Definition 1 with regret level $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$ and confidence level $1-\alpha$.
@@
To lighten notations, denote by $r$ the actual maximum regret $\operatorname{\textbf{rgt}}^{\textup{max}}_{\hat{\theta}}(v_{\textup{test}};b_{\textup{test}})$ and by $\hat{r}$ the predicted maximum regret ${r}_{\hat{\theta}_{r}}^{\textup{max}}(b_{\textup{test}})$. With these notations in place, the probability of interest is as follows:
$\mathbb{P}\left[(r>\operatorname{\ell_{\operatorname{\textbf{rgt}}}})\>\cap\>(\hat{\mathcal{T}}_{\alpha}(b_{\textup{test}},\operatorname{\ell_{\operatorname{\textbf{rgt}}}})\neq\emptyset)\right]$
@@
= $=\>1-\mathbb{P}\left[r-\hat{r}\leq\hat{q}_{\alpha}\right]$ complement event
 $\leq\>1-(1-\alpha)$
@@
where the last inequality holds due [Romano et al., 2019, Lemma 2]. ∎
4 Experiments
In this section, we empirically evaluate our proposal on auctions with configurations of (bidders $\times$ items) sets of size $2\times 2$, $2\times 3$, and $3\times 5$. The training process is based on the RegretNet implementation by Dütting et al. [2019] (PyTorch version by Stein et al. [2023]), which is used as the data-driven auction model $f_{\hat{\theta}}$. Throughout the experiments, we refer to our conformal method as ConfRegretNet. That is, we apply the acceptance rule $\hat{\mathcal{T}}_{\alpha}$ to the output of the auction model ${f}_{\hat{\theta}}(b)=[{g}_{\hat{\theta}}(b),{p}_{\hat{\theta}}(b)]$. We fit our regret estimation model ${r}_{\hat{\theta}_{r}}$ within the shared backbone architecture setup, illustrated in 3. For further implementation details, please refer to Section A in the Appendix.
The performance of our ConfRegretNet method is evaluated in terms of revenue, empirical regret, and maximum regret. Table 1 summarizes the results for different auction configurations. Following that table, we can see our method achieves competitive revenue compared to RegretNet across different auction settings while controlling the requested levels of maximum regret. Observe how the maximum regret values obtained by RegretNet are uncontrolled.
(a) Auction setup 2x2
(b) Auction setup 2x3
(c) Auction setup 3x5
Figure 5: The distribution of regret values corresponding to RegretNet and ConfRegretNet across different auction settings: $2\times 2$, $2\times 3$, and $3\times 5$. The test and the calibration sets are of size 1,000 examples each. The red dashed line indicates the requested regret level $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$ of each setup.
Figure 6: Scatter plot of the actual regret $\operatorname{\textbf{rgt}}_{\hat{\theta}}(v;b)$ and the predicted regret ${r}_{\hat{\theta}_{r}}(b)$ of ConfRegretNet, evaluated on test data of size $1,000$; the auction setup is $2\times 2$.
Table 1 also shows that our acceptance rule tends to accept most of the auctions at test time. Further, Figure 5 illustrates that (i) the auctions that are rejected by our method are indeed those that exceed the specified regret level; and (ii) the regret values of the accepted auctions fall below the maximal regret level $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$. This implies that our approach not only controls the regret level but also tends to accept auctions that should be accepted.
Table 2 summarizes the performance of various classical, transformer-based, and deep learning-based auction mechanisms in the $2\times 2$ setting. Among the strictly strategy-proof methods, AMenuNet achieves a higher revenue (0.43) than the classical VCG (0.28) and Item-Myerson (0.42). However, better revenues are achieved by the neural network models—RegretFormer (0.94), RegretNet (0.93), ALGNet (0.88), and EquivariantNet (0.87)—all of which surpass the classical approaches. Notably, ConfRegretNet attains revenue (0.92) comparable to these high-performing models while offering a statistical strategy-proofness guarantee. This demonstrates the practical trade-off between guaranteeing zero-regret and maximizing revenue, highlighting ConfRegretNet as a compelling option for real-world auction applications where both factors matter.
Table 2: Comparison of average revenue and strategy-proofness of various auction models in the $2\times 2$ auction setting. The table includes our ConfRegretNet compared to classical mechanisms (VCG, Item-Myerson), transformer-based models (AMenuNet, RegretFormer), and other deep learning-based approaches (RegretNet, EquivariantNet, ALGNet). Strategy-proofness is indicated as either Yes, No, or Statistically. The results presented in this table are as reported by the authors for the same setup.
Auction Model Revenue Strategy-Proof
VCG [Vickrey, 1961] 0.28 Yes
Item-Myerson [Myerson, 1981] 0.42 Yes
AMenuNet [Duan et al., 2023] 0.43 Yes
RegretNet [Dütting et al., 2019] 0.93 No
EquivariantNet [Rahme et al., 2021a] 0.87 No
ALGNet [Rahme et al., 2021b] 0.88 No
RegretFormer [Ivanov et al., 2022] 0.94 No
ConfRegretNet [ours] 0.92 Statistically
Table 3: Summary of the average revenue and empirical regret (with standard deviations in parentheses) for RegretNet, ConfRegretNet-shared (3), and ConfRegretNet-separate (4) in the black-box settings described in Section 3.1. Each model’s requested and actual maximum regret is reported, along with the percentage of auctions accepted under the requested regret threshold. Test and calibration sets each contain 10,000 examples.
Auction Model Auction Setup Bidders × Items Revenue Empirical Regret Requested Max Regret Max Regret Acceptance Auctions
RegretNet $2\times 2$ 0.93 (0.33) 0.007 (0.006) - 0.039 -
ConfRegretNet-shared (3) $2\times 2$ 0.92 (0.32) 0.004 (0.003) 0.025 0.020 89.23%
ConfRegretNet-separate (4) $2\times 2$ 0.92 (0.31) 0.004 (0.005) 0.025 0.023 93.7%
Figure 7: The effect of the maximal allowed regret level $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$ on the number of auctions rejected by the calibrated acceptance rule $\hat{\mathcal{T}}_{\alpha}(b,\operatorname{\ell_{\operatorname{\textbf{rgt}}}})$. Results are evaluated on a test set of size 400; the auction setup is $2\times 2$.
5 Discussion
In this paper, we introduced the notion of statistical strategy-proof auctions, and formulated a data-driven acceptance rule that selectively accepts auctions, at test time, whose regret is guaranteed to fall below a user-specified maximum regret value with high probability. The formulation of our data-driven acceptance rule is achieved based on predictions obtained by our novel regret estimation model, which is then calibrated by building on ideas from the conformal prediction literature.
Numerical experiments showed that while the average regret of the vanilla RegretNet model is low, its maximum regret can be relatively high. By applying our proposed acceptance rule, the regret of accepted auctions consistently remained within the prescribed maximum regret level. Notably, the results demonstrated the effectiveness of our approach in filtering out high-regret auctions while retaining those with low regret. The experiments also demonstrated that our approach achieved notably higher revenue compared to baseline methods that enforce strictly strategy-proof auctions with zero regret, highlighting the practical relevance of our statistical approach.
Limitations and future directions
A natural limitation of our method is the i.i.d. assumption we make, which might be violated in practice. This may happen due to changes in the bidding and valuation distribution profiles, or due to imperfect calibration data that include noisy or inaccurate valuations, for instance. As a future research direction, it would be important to quantify the effect of violation of this assumption on the resulting maximal regret level, possibly by taking inspiration from Barber et al. [2023]. It would be also intriguing to design calibration frameworks that are robust to such distribution shifts in the data.
Another worthwhile direction for future work is to extend our framework to simultaneously control the test-time revenue and regret. We believe this could be done by borrowing ideas from [Angelopoulos et al., 2021], which presents a general calibration framework for controlling multiple risks. Such an approach would enable the auctioneer to ensure, with high probability, that the calibrated auction mechanism maintains low regret while also keeping the revenue of the acceptance auctions above a specified threshold. This multi-objective control is particularly relevant for large-scale, multi-item auctions, where revenue could potentially fall dramatically due to a large number of items.
Acknowledgements.
Supplementary Material
Appendix A Architectural and Training Details

### v5

@@ 3 Proposed Method
Definition 3.1 (Statistically Strategy-Proof Auctions).
@@
In words, an auction outcome that satisfies Definition 3.1 is guaranteed to have low maximum regret with high probability, where the maximal regret value $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$ and the confidence level $1-\alpha$ are pre-defined by the auctioneer. Since we can control the maximal regret value with high probability, we refer to it as a statistically strategy-proof auction mechanism. Notably, this statistical notion of strategy-proofness does not imply zero regret at test time.
@@ 3.1 Regret Estimation
We propose two different approaches to design the regret estimation model ${r}_{{\hat{\theta}}_{r}}(b)$. In the shared backbone setup, the regret model is integrated into the existing allocation and payment backbone, providing a unified structure. The advantage of this approach is that it leverages the backbone features of the auction model to better predict the regret. However, this setup requires access to the auction model parameters and modifies the training procedure. In the second black-box approach, the auction model is treated as a black-box, avoiding the need to access the auction model’s internal parameters. Illustrations of both setups are provided in Appendix E.
@@ 3.2 Model Calibration
In essence, we show how to form a corrected threshold on the output ${r}^{\textup{max}}_{\hat{\theta}_{r}}(b_{\text{test}})$, ensuring the regret of the auction outcome obtained from the auction model does not exceed the requested maximal regret level $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$ with probability $1-\alpha$, resulting in a calibrated auction acceptance rule function. Given the auction model $^{f}_{\hat{\theta}}$, a calibration set $\mathcal{D}_{\textup{cal}}$, a requested regret level $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$, and a control level $\alpha$, we define the auction acceptance rule function $\hat{\mathcal{T}}_{\alpha}(b_{\textup{test}},\operatorname{\ell_{\operatorname{\textbf{rgt}}}})$ for a test bid vector $b_{\text{test}}$ as follows:
$\hat{\mathcal{T}}_{\alpha}(b_{\text{test}},\operatorname{\ell_{\operatorname{\textbf{rgt}}}})=\begin{cases}{f}_{\hat{\theta}}(b_{\text{test}}),&{r}^{\textup{max}}_{\hat{\theta}_{r}}(b_{\text{test}})\leq\operatorname{\ell_{\operatorname{\textbf{rgt}}}}-\hat{q}_{\alpha},\\ \;\;\;A_{0},&\text{otherwise}.\end{cases}$ (8)
The function $\hat{\mathcal{T}}_{\alpha}(b_{\text{test}},\operatorname{\ell_{\operatorname{\textbf{rgt}}}})$ from above is determined by the estimated maximum regret ${r}^{\textup{max}}_{\hat{\theta}_{r}}(b_{\text{test}})$ and the calibration threshold $\hat{q}_{\alpha}$, which is derived using the calibration set $\mathcal{D}_{\textup{cal}}$. The output of the function $\hat{\mathcal{T}}_{\alpha}(b_{\text{test}},\operatorname{\ell_{\operatorname{\textbf{rgt}}}})$ is the prediction of the auction model ${f}_{\hat{\theta}}({b_{\text{test}}})$ if ${r}^{\textup{max}}_{\hat{\theta}_{r}}(b_{\text{test}})$ is lower than the difference between the requested regret level $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$, and the correction factor $\hat{q}_{\alpha}$ derived from our calibration method. Otherwise, the function $\hat{\mathcal{T}}_{\alpha}$ returns no-allocation rule $A_{0}$, indicating that the predicted auction does not meet the desired regret level. This acceptance rule ensures that the auction mechanism only operates within the predefined regret bounds.
@@
1:Input: $^{f}_{\hat{\theta}}$ – pre-trained auction model; $r_{\hat{\theta}_{r}}$ – pre-trained regret model; $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$ – max regret level; $(1-\alpha)$ – requested confidence level; $\mathcal{D}_{\textup{cal}}$ – holdout calibration set.
@@
4:  Compute the estimated max regret ${r}_{\hat{\theta}_{r}}^{\textup{max}}(b_{k})$.
5:  Compute the actual max regret $\operatorname{\textbf{rgt}}_{\hat{\theta}}^{\textup{max}}(v_{k};b_{k})$.
6:  Calculate $e_{k}=\operatorname{\textbf{rgt}}_{\hat{\theta}}^{\textup{max}}(v_{k};b_{k})-{r}_{\hat{\theta}_{r}}^{\textup{max}}(b_{k})$.
@@
Table 1: Summary of the average revenue and empirical regret, with their standard deviations in parentheses, for both RegretNet and SPACE across $2\times 2$, $2\times 3$, and $3\times 5$ auction settings. The test and the calibration sets are of size 1,000 examples. The ‘Requested Max Regret Level‘ column indicates the requested regret used in SPACE to control the regret at $99\%$ confidence ($\alpha=0.01$). The ‘Accepted Auctions‘ column shows the percentage of accepted auctions by SPACE based on the regret threshold. the ‘Max Regret‘ column refers the max regret obtained at test time.
Auction Model
Auction Setup
Bidders × Items
Revenue
Empirical
Regret
Requested
Max Regret
Max
Regret
Acceptance
Auctions
RegretNet $2\times 2$ 0.93 (0.33) 0.007 (0.007) 0.025 0.032 100%
SPACE [ours] $2\times 2$ 0.90 (0.31) 0.005 (0.004) 0.025 0.024 93.9%
RegretNet $2\times 3$ 1.38 (0.37) 0.007 (0.008) 0.035 0.045 100%
SPACE [ours] $2\times 3$ 1.34 (0.38) 0.006 (0.005) 0.035 0.033 95.2%
RegretNet $3\times 5$ 2.91 (0.40) 0.012 (0.012) 0.055 0.073 100%
SPACE [ours] $3\times 5$ 2.60 (0.39) 0.010 (0.009) 0.055 0.054 97.7%
We now prove that any test auction accepted by the rule $\hat{\mathcal{T}}_{\alpha}$ satisfies Definition 3.1. Simply put, this result guarentees that by setting $\alpha=0.01$ and specifying a maximal regret level of $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}=0.025$, the actual regret of new auctions accepted by $\hat{\mathcal{T}}_{\alpha}$ will not exceed 0.025 with probability 99%.
@@
Suppose the calibration set $\mathcal{D}_{\textup{cal}}=\{(b_{k},v_{k})\}_{k=1}^{K}$ and a new auction $(b_{\textup{test}},v_{\textup{test}})$ are i.i.d samples from $P_{b,v}$. Let $\hat{\mathcal{T}}_{\alpha}$ be the acceptance rule of a given auction model $^{f}_{\hat{\theta}}$ and regret estimator $r_{\hat{\theta}_{r}}$, obtained from Algorithm 1. Then, the auction ${f}_{{\hat{\theta}}}(b_{\textup{test}})$ accepted by the rule $\hat{\mathcal{T}}_{\alpha}$ in (8) satisfies Definition 3.1 with regret level $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$ and confidence level $1-\alpha$.
4 Experiments
In this section, we empirically evaluate our proposal on auctions with configurations of (bidders $\times$ items) sets of size $2\times 2$, $2\times 3$, and $3\times 5$. The training process is based on the RegretNet implementation by Dütting et al. [11] (PyTorch version by [35]), which is used as the data-driven auction model $f_{\hat{\theta}}$. Throughout the experiments, we refer to our conformal method as SPACE . That is, we apply the acceptance rule $\hat{\mathcal{T}}_{\alpha}$ to the output of the auction model ${f}_{\hat{\theta}}(b)=[{g}_{\hat{\theta}}(b),{p}_{\hat{\theta}}(b)]$. We fit our regret estimation model ${r}_{\hat{\theta}_{r}}$ within the shared backbone architecture setup. For further implementation details, please refer to Appendix D.
The performance of our SPACE method is evaluated in terms of revenue, empirical regret, and maximum regret. Table 1 summarizes the results for different auction configurations. Following that table, we can see our method achieves competitive revenue compared to RegretNet across different auction settings while controlling the requested levels of maximum regret. Observe how the maximum regret values obtained by RegretNet are uncontrolled.
Figure 4: Distribution of regret values for RegretNet and SPACE in the $2\times 3$ auction setting. The test and calibration sets each contain 1,000 examples. The red dashed line indicates the requested regret level $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$.
Figure 5: Boxplot of the true regret values for SPACE in the $2\times 3$ auction setting. The test and calibration sets each contain 1,000 examples. The red dashed line indicates the requested regret level $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$.
Table 1 also shows that our acceptance rule tends to accept most auctions at test time. Further, Figure 4 illustrates that (i) the auctions rejected by our method are those that exceed the specified regret level; and (ii) the regret values of the accepted auctions fall below the maximal regret level $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$. Consistently, Fig. 5 shows rejected auctions are concentrated above $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$, while accepted auctions are concentrated below it. Taken together, these results imply that SPACE controls regret, rejects high-regret auctions, and tends not to reject low-regret auctions—i.e., it accepts the auctions that should be accepted.
Table 2: Comparison of average revenue and strategy-proofness of various auction models in the $2\times 2$ auction setting. The table includes our SPACE compared to classical mechanisms (VCG, Item-Myerson), transformer-based models (AMenuNet, RegretFormer), and other deep learning-based approaches (RegretNet, EquivariantNet, ALGNet). Strategy-proofness is indicated as either Yes, No, or Statistically. The results presented in this table are as reported by the authors for the same setup.
Auction Model
Revenue
Strategy-Proof
VCG [36] 0.28 Yes
Item-Myerson [23] 0.42 Yes
AMenuNet [9] 0.43 Yes
RegretNet [11] 0.93 No
EquivariantNet [27] 0.87 No
ALGNet [28] 0.88 No
RegretFormer [14] 0.94 No
SPACE  (ours) 0.92 Statistically
Figure 6: Effect of $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$ on the number of auctions rejected by the calibrated acceptance rule $\hat{\mathcal{T}}_{\alpha}(b,\operatorname{\ell_{\operatorname{\textbf{rgt}}}})$ in the $2\times 2$ auction setup. Evaluated on a test set of 400 examples.
5 Discussion
In this paper, we introduced the notion of statistical strategy-proof auctions, and formulated a data-driven acceptance rule that selectively accepts auctions, at test time, whose regret is guaranteed to fall below a user-specified maximum regret value with high probability. The formulation of our data-driven acceptance rule is achieved based on predictions obtained by our novel regret estimation model, which is then calibrated by building on ideas from the conformal prediction literature. Numerical experiments showed that while the average regret of the vanilla RegretNet model is low, its maximum regret can be relatively high. By applying our proposed acceptance rule, the regret of accepted auctions consistently remained within the prescribed maximum regret level. Notably, the results demonstrated the effectiveness of our approach in filtering out high-regret auctions while retaining those with low regret. The experiments also demonstrated that our approach achieved notably higher revenue compared to baseline methods that enforce strictly strategy-proof auctions with zero regret, highlighting the practical relevance of our statistical approach.
A natural limitation of our method is the i.i.d. assumption we make, which might be violated in practice. This may happen due to changes in the bidding and valuation distribution profiles, or due to imperfect calibration data that include noisy or inaccurate valuations, for instance. As a future research direction, it would be important to quantify the effect of violation of this assumption on the resulting maximal regret level, possibly by taking inspiration from [4]. It would be also intriguing to design calibration frameworks that are robust to such distribution shifts in the data.
Another worthwhile direction for future work is to extend our framework to simultaneously control the test-time revenue and regret. We believe this could be done by borrowing ideas from [2], which presents a general calibration framework for controlling multiple risks. Such an approach would enable the auctioneer to ensure, with high probability, that the calibrated auction mechanism maintains low regret while also keeping the revenue of the acceptance auctions above a specified threshold. This multi-objective control is particularly relevant for large-scale, multi-item auctions, where revenue could potentially fall dramatically due to a large number of items.
Acknowledgements
Appendix A Additional Experiments
This appendix provides additional experimental results and diagnostics that complement the main text. We report extended quantitative metrics—coverage of the statistical truthfulness constraint, realized regret relative to the requested level $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$, and revenue—alongside robustness checks to calibration set size, model capacity, and distribution shift, and ablations of the components of SPACE . Unless stated otherwise, the data generation process, training protocol, and hyperparameters are identical to those in Sec. 4.
Table 3 mirrors Table 1 but evaluates on larger calibration and test sets (10,000 each). The pattern is unchanged: SPACE achieves revenue close to RegretNet while markedly lowering empirical regret and keeping test-time max regret at or below the requested level $\operatorname{\ell_{\operatorname{\textbf{rgt}}}}$ in all settings. By contrast, RegretNet exceeds the requested max-regret level in each case. The increased sample size tightens the estimates and demonstrates that SPACE scales to large calibration and test sets while preserving coverage and revenue.
Table 3: Summary of the average revenue and empirical regret, with their standard deviations in parentheses, for both RegretNet and SPACE across $2\times 2$, $2\times 3$, and $3\times 5$ auction settings. The test and the calibration sets are of size 10,000 examples. The ‘Requested Max Regret Level‘ column indicates the requested regret used in SPACE to control the regret at $99\%$ confidence ($\alpha=0.01$). The ‘Accepted Auctions‘ column shows the percentage of accepted auctions by SPACE based on the regret threshold. the ‘Max Regret‘ column refers the max regret obtained at test time.
Auction Model
Auction Setup
Bidders × Items
Revenue
Empirical
Regret
Requested
Max Regret
Max
Regret
Acceptance
Auctions
RegretNet $2\times 2$ 0.93 (0.33) 0.007 (0.006) 0.025 0.039 100%
SPACE [ours] $2\times 2$ 0.92 (0.32) 0.004 (0.003) 0.025 0.020 89.23%
RegretNet $2\times 3$ 1.41 (0.38) 0.011 (0.014) 0.035 0.164 100%
SPACE [ours] $2\times 3$ 1.38 (0.39) 0.007 (0.005) 0.035 0.032 94.4%
RegretNet $3\times 5$ 2.93 (0.40) 0.023 (0.012) 0.055 0.121 100%
SPACE [ours] $3\times 5$ 2.67 (0.38) 0.009 (0.007) 0.055 0.055 93.4%
Table 4: Summary of the average revenue and empirical regret (with standard deviations in parentheses) for RegretNet, SPACE -shared (8(a)), and SPACE -separate (8(b)) in the black-box settings described in Section 3.1. Each model’s requested and actual maximum regret is reported, along with the percentage of auctions accepted under the requested regret threshold. Test and calibration sets each contain 10,000 examples.
Auction Model
Auction Setup
Bidders × Items
Revenue
Empirical
Regret
Requested
Max Regret
Max
Regret
Acceptance
Auctions
RegretNet $2\times 2$ 0.93 (0.33) 0.007 (0.006) 0.025 0.039 100%
SPACE -shared (8(a)) $2\times 2$ 0.92 (0.32) 0.004 (0.003) 0.025 0.020 89.23%
SPACE -separate (8(b)) $2\times 2$ 0.92 (0.31) 0.004 (0.005) 0.025 0.023 93.7%
Table 5 mirrors Table 2 and expands it with additional bidder–item configurations, showing that SPACE scales cleanly with both the number of bidders and items. Across all settings, SPACE attains revenue that is competitive with the strongest learning-based baselines. Unlike these high-revenue DNN baselines, which are not strategy-proof, SPACE preserves statistical strategy-proofness at test time. Classical truthful mechanisms provide strict strategy-proofness but generally trail the top non-truthful models in revenue for larger setups. Overall, the extended results confirm a favorable revenue–truthfulness trade-off: SPACE maintains near–state-of-the-art revenue while delivering test-time guarantees and scaling gracefully as the auction size grows.
Table 5: Comparison of average revenue and strategy-proofness of various auction models in different auction settings. The table includes our SPACE compared to classical mechanisms (VCG, Item-Myerson), transformer-based models (AMenuNet, RegretFormer), and other deep learning-based approaches (RegretNet, EquivariantNet, ALGNet). Strategy-proofness is indicated as either Yes, No, or Statistically. The results presented in this table are as reported by the authors for the same setups.
Auction Model
1x2
2x2
2x3
3x10
Strategy-Proof
VCG [36] 0.00 0.28 1.00 5.02 Yes
Item-Myerson [23] 0.14 0.42 1.25 5.31 Yes
AMenuNet [9] 0.17 0.43 1.30 5.58 Yes
RegretNet [11] 0.57 0.93 1.40 5.59 No
EquivariantNet [27] 0.58 0.87 1.42 5.74 No
ALGNet [28] 0.55 0.88 1.44 5.52 No
RegretFormer [14] 0.57 0.94 1.48 6.12 No
SPACE  (ours) 0.57 0.92 1.39 5.52 Statistically
Figure 7: Scatter plot of the actual regret $\operatorname{\textbf{rgt}}_{\hat{\theta}}(v;b)$ and the predicted regret ${r}_{\hat{\theta}_{r}}(b)$ of SPACE , evaluated on test data of size $1,000$; the auction setup is $2\times 2$.
Appendix B Mathematical Proofs
@@
of Theorem 1. To lighten notations, denote by $r$ the actual maximum regret $\operatorname{\textbf{rgt}}^{\textup{max}}_{\hat{\theta}}(v_{\textup{test}};b_{\textup{test}})$ and by $\hat{r}$ the predicted maximum regret ${r}_{\hat{\theta}_{r}}^{\textup{max}}(b_{\textup{test}})$. With these notations in place, the probability of interest is as follows:
$\mathbb{P}\left[(r>\operatorname{\ell_{\operatorname{\textbf{rgt}}}})\>\cap\>(\hat{\mathcal{T}}_{\alpha}(b_{\textup{test}},\operatorname{\ell_{\operatorname{\textbf{rgt}}}})\neq A_{0})\right]$
@@
= $=1-\mathbb{P}\left[r-\hat{r}\leq\hat{q}_{\alpha}\right]$ complement event
 $\geq 1-(1-\alpha)$
@@
where the last inequality holds due [30, Lemma 2]. ∎
Appendix C Experiment setup and environment
The experiments were conducted on a high-performance computing setup consisting of an Ubuntu 20.04.6 LTS operating system, powered by 96 Intel(R) Xeon(R) Gold CPUs running at 2.40 GHz. The machine was equipped with 16 Nvidia A40 GPUs and 512 GB of RAM. The software environment was based on PyTorch 2.1, running under Python 3.11.5.
Appendix D Architectural and Training Details